# Critical spin fluctuations across the superconducting dome in La2-xSrxCuO4

### v1

@@ Abstract
Overdoped cuprate superconductors are strange metals above their superconducting transition temperature. In such materials, the electrical resistivity has a strong linear dependence on temperature ($T$) and electrical current is not carried by electron quasiparticles as in conventional metals. Here we demonstrate that the strange metal behavior co-exists with strongly temperature-dependent critical spin fluctuations showing dynamical scaling across the cuprate phase diagram. Our neutron scattering observations and the strange metal behavior are consistent with a spin density wave quantum phase transition in a metal with spatial disorder in the tuning parameter. Numerical computations on a Yukawa-Sachdev-Ye-Kitaev model yield an extended ‘Griffiths phase’ with scaling properties in agreement with observations, establishing that low-energy spin excitations and spatial disorder are central to the strange metal behavior.
@@
Understanding high-temperature superconductivity in layered cuprates has posed significant challenges to the theory of condensed matter. In particular, the unusual metallic state of cuprates above their critical temperature, $T_{c}$, remains poorly understood. This state is crucial, as superconductivity emerges from it. The cuprate phase diagram is characterized by the presence of various competing orders such as antiferromagnetism (AF), charge density wave (CDW) order, and the pseudogap state in the underdoped region of the phase diagram Keimer et al. (2015). The overdoped region is largely free from these competing phases. Nevertheless, the normal state ($T>T_{c}$) shows “strange metal” (SM) behavior Phillips et al. (2023) where the resistivity, $\rho$, is proportional to temperature, $T$, with a large proportionally constant. The resistivity of a metal is usually interpreted using the Drude model in which $\rho\propto\tau_{\text{tr}}^{-1}$, where $\tau_{\text{tr}}^{-1}$ is the relaxation rate of the electrons transporting current. For SMs, we find that $\tau_{\text{tr}}^{-1}\approx k_{B}T/\hbar$, the Planckian dissipative rate Zaanen (2004). This is in stark contrast to the $\rho\propto T^{2}$ and $\tau_{\text{tr}}^{-1}\propto T^{2}$ expected to describe metals in Fermi liquid theory at low temperature. The strange behavior Hartnoll and Mackenzie (2022) in resistivity is accompanied by an anomalous specific heat contribution $\propto T\ln(1/T)$ in contrast to the liquid Fermi form $\propto T$.
The existence of the Planckian transport relaxation rate $\tau_{\text{tr}}^{-1}(T)$ over a wide temperature range in cuprates is challenging for theory to explain. The picture in simple metals is based on the scattering of the fermion quasiparticles near the Fermi surface off bosons (e.g. phonons). The total momentum of the system is then degraded via impurities or umklapp scattering. There are two issues with is picture with regards to cuprates Patel et al. (2023). First, linear or nearly-linear resistivity exists over a wide range of doping so the relevant excitations would also be present over this range. Second, the degradation of momentum mentioned above cannot be accounted for by umklapp scattering. Thus other mechanisms are required.
@@
Here we show that nearly-critical low-energy collective spin fluctuations are integral to the strange metal behavior in the cuprates and exist across the superconducting dome. We investigate the overdoped region a cuprate superconductor where the SM behavior is clearest Cooper et al. (2009). The material we have chosen is La2-xSrxCuO4 (LSCO). This system can be doped over a wide range through the overdoped region of the phase diagram (Fig. 1A) and has a relatively low $T_{c}$ so that the SM normal state is present over a wide range of temperature.
@@
Figure 1: Spin fluctuations in La2-xSrxCuO4 (x=0.22). (A) Schematic phase diagram of LSCO showing the critical spin fluctuations observed here. (B) Schematic of spin excitations in LSCO($x=0.22$) based on (Refs. Lipscombe et al. (2007); Robarts et al. (2019)) as represented by imaginary part of the susceptibility $\chi^{\prime\prime}(\mathbf{Q},\omega)$. Reciprocal space is labeled as $\mathbf{Q}=H\mathbf{a}^{\star}+K\mathbf{b}^{\star}+L\mathbf{c}^{\star}\equiv(H,K)$. Gray region is broad high energy spin excitations Lipscombe et al. (2007); Robarts et al. (2019). Pink regions are the low-energy excitations studied here. (C-F) Slices of $\chi^{\prime\prime}(\mathbf{Q},\omega)$ for $\hbar\omega=4$ meV, $L\in[-1,1]$ and various temperatures (seeSupplementary Sup for details). White dashed lines indicate the range over which slices are integrated to produce cuts in Fig. 2.
@@
Figure 2: The temperature dependence to the low-energy spin fluctuations in La2-xSrxCuO4 (x=0.22). (A) The scattering intensity versus wavevector $\mathbf{Q}$ for different temperatures (bottom to top: 26, 40, 80, 155, 300 K) for energy transfers $\hbar\omega=4,8$ meV. The cuts are through the incommensurate $\mathbf{Q}_{\delta}$ positions as shown by the dashed lines in Fig. 1D. (B) $E-\mathbf{Q}$ map of the magnetic response function $\chi^{\prime\prime}(\mathbf{Q},\omega)$ for $T=26,80$ K showing its evolution with temperature. The trajectory of $\mathbf{Q}$ is the same as that in (A).
@@
Figure 3: Scaling behavior of the dynamic magnetic susceptibility and correlation length in La2-xSrxCuO4 (A) Temperature dependence of $\chi^{\prime\prime}(\mathbf{Q},\omega)$ at ordering wavevector $\mathbf{Q}_{\delta}$. Lines are fits to Eqn. 3. Inset shows the $T$-dependence of the spin relaxation rate $\Gamma_{\delta}$ and real part of the susceptibility $\chi^{\prime}(\mathbf{Q}_{\delta})$ obtained when data are modeled with Eqns. 3-4. (B) $\omega/T$ scaling plot of the $x=0.22$ data in (A) using Eqn. 2 and the procedure described in Sup . The inset shows how the quality of the collapse Sup varies with exponent $\alpha$. (C) Scaling plot for underdoped LSCO $x=0.14$ data from Ref. Aeppli et al. (1997). This data shows a similar collapse to the $x=0.22$ data in (B). (D) The width in $\mathbf{Q}$ of $\chi^{\prime\prime}(\mathbf{Q,\omega})$ is denoted by the inverse dynamic correlation length $\kappa(\omega)$. Here we plot $\kappa(\omega)$ versus temperature for $\hbar\omega=1.25,2.25,4,8$ meV. (E) Data in (D) can be collapsed onto a single trend line with suitable choice of $z$ and $\omega/T$ scaling. (F) Data in (D) can be scaled onto a single line using (7) allowing the dynamic critical exponent $z$ to be determined. Dashed line is Eqn. S15 Sup .
@@
We used inelastic neutron scattering to map about the $\mathbf{Q}$-$\omega$ dependence of the low-energy spin fluctuation in single crystals of LSCO(x$=0.22$) in the normal state for temperatures $T_{c}\leq T\leq 300$ K. Measurement were carried out using the LET spectrometer of the ISIS Pulsed Neutron and Muon Source Sup . Fig. 1C-F show constant energy maps of the dynamic susceptibility $\chi^{\prime\prime}(\mathbf{Q},\omega)$ at $\hbar\omega=4$ meV. For $T=26$ K, the four-peaks around $\mathbf{Q}=(1/2,1/2)$ can be clearly seen in Fig. 1C. Cuts through the peaks along the $\mathbf{Q}=(1/2-\delta/2+\xi,1/2-\delta/2+\xi)$ line (see dashed lines in Fig. 1D) are shown in Fig. 2A. As the temperature increases, the peaks can be seen to weaken and broaden in $\mathbf{Q}$. The measured intensity $S(\mathbf{Q},\omega)$ can be converted to the the magnetic response function $\chi^{\prime\prime}(\mathbf{Q},\omega)$ using the fluctuation dissipation theorem $S(\mathbf{Q},\omega)=(1/\pi)\chi^{\prime\prime}(\mathbf{Q},\omega)[1-\exp(-\hbar\omega/k_{B}T)]^{-1}$ Sup , the results shown in Fig. 2B and Fig. 3A. Our data (see Fig. 3A) show that the dynamic spin susceptibility near $\mathbf{Q}_{\delta}$ is strongly temperature dependent for this over-doped cuprate. The behavior is reminiscent of the approach to a quantum critical point as $T\rightarrow 0$.
At a quantum critical point, the dynamical susceptibility is expected to show the scaling behavior Sup ; Sachdev and Ye (1992)
@@
where $\xi\propto T^{-1/z}$ is a correlation length, $z$ is the dynamic critical exponent, $q=\left|\mathbf{Q}-\mathbf{Q}_{\delta}\right|$ is wavevector relative to ordering wavevector, and $\Phi$ is a universal complex function of both arguments. For $q=0$, the imaginary part of the dynamic susceptibility (measured here) scales Sup as
@@
where $\alpha=\gamma/\nu z$, $\phi_{1}(x)=\Phi(0,x)$ is a scaling function and $\gamma$ and $\nu$ are the susceptibility and correlation function temperature exponents.
In Fig. 3B we show that our data can be collapsed onto a single trend with a suitable choice of $\alpha$, where $\phi_{1}(x)$ can be approximated by a simple Lorentzian $\phi_{1}(x)\propto ax/(a^{2}+x^{2})$ with $a=2.9\pm 0.3$ and $\alpha=0.32\pm 0.10$. This small value of $\alpha$ excludes the familiar paramagnon theory which has $\alpha=1$ Millis et al. (1990) (which also does not have an extended regime of criticality). Our data is consistent with the susceptibility having a low-frequency component that varies as
@@
where the real part of the susceptibility $\chi^{\prime}(\mathbf{Q}_{\delta})\propto T^{-\alpha}$ and spin relaxation rate varies as
@@
as shown inset of Fig. 3A. Thus our data is consistent with $\Gamma_{\delta}$ tracking the Planckian relaxation rate $\tau^{-1}_{\text{tr}}\propto T$ as observed in transport measurements suggesting that they are related.
We can gain more information about the criticality of the spin fluctuations from the $\mathbf{Q}$-dependence of $\chi^{\prime\prime}(\mathbf{Q},\omega)$. From Eqn. 3 we see that $\chi^{\prime\prime}(q,\omega)\propto\omega$ as $\omega\rightarrow 0$ so there will be peaks in $\chi^{\prime\prime}(\mathbf{Q},\omega)/\omega$ at the lowest energies. It can be shown from the scaling of the susceptibility Sup that
$\left.\frac{\chi^{\prime\prime}(q,\omega)}{\omega}\right|_{\omega\rightarrow 0}\propto T^{-\left(\alpha+1\right)}\phi_{2}(\xi q).$ (5)
The function $\phi_{2}(\xi q)$ is peaked at the ordering wavevector $q=0$ with width $\xi^{-1}$. We can determine $z$ from the temperature dependence of $\xi^{-1}$.
In order to fit our data, we use a form previously used Aeppli et al. (1997); Zhu et al. (2023) to describe the low-energy response of the cuprates:
$\chi^{\prime\prime}(\mathbf{Q},\omega)$ $=\chi^{\prime\prime}(\mathbf{Q}_{\delta},\omega)\frac{\kappa^{4}(\omega)}{[\kappa^{2}(\omega)+R(\mathbf{Q})]^{2}},$ (6)
where $R(\mathbf{Q})$ is a function with zeros at the $\mathbf{Q}_{\delta}$ positions and the property $R(\mathbf{Q}_{\delta})=|\mathbf{q}|^{2}$ for $\mathbf{q}\rightarrow 0$, the second factor represents $\phi_{2}(\xi q)$ in the limit $\omega\rightarrow 0$ and we can make the association $\kappa(\omega=0)\equiv\xi^{-1}$ Sup . Example fits to Eqn. 6 are shown in Fig. 2. The resulting fitted values of $\kappa(\omega)$ ($\mathbf{Q}$-width of response) are shown in Fig. 3D,E,F. In Fig. 3D we see that $\kappa(\omega)$ increases with $\omega$ and $T$. We posit that the effects of $\omega$ and $T$ add in quadrature and this suggests the form
@@
where $\kappa(\omega=0)=\xi^{-1}\propto T^{1/z}$ and $r$ is the ratio of the temperature and energy contributions which we set equal to one. We find that Eqn. 7 can be used to collapse the experimentally determined $\kappa(\omega,T)$ onto a single solid line as shown in Fig. 3E,F with $z=1.83\pm 0.35$. In the Supplementary Sup we show that our data are also consistent with the system being tuned slightly away from quantum criticality (dashed line in Fig. 3F).

### v2

@@ Abstract
Overdoped cuprate superconductors are strange metals above their superconducting transition temperature. In such materials, the electrical resistivity has a strong linear dependence on temperature ($T$) and electrical current is not carried by electron quasiparticles as in conventional metals. Here we demonstrate that the strange metal behaviour co-exists with strongly temperature-dependent critical spin fluctuations showing dynamical scaling across the cuprate phase diagram. Our neutron scattering observations and the strange metal behaviour are consistent with a spin density wave quantum phase transition in a metal with spatial disorder in the tuning parameter. Numerical computations using a theory of spin density waves in a disordered metal yield an extended ‘Griffiths phase’ with scaling properties in agreement with experimental observations. Thus we establish that low-energy spin excitations and spatial disorder are central to the strange metal behaviour.
@@
Understanding high-temperature superconductivity in layered cuprates has posed significant challenges to the theory of condensed matter. In particular, the unusual metallic state of cuprates above their critical temperature, $T_{c}$, remains poorly understood. This state is crucial, as superconductivity emerges from it. The underdoped region of the cuprate phase diagram is characterized by the presence of various competing orders such as antiferromagnetism (AF), charge density wave (CDW) order, and the pseudogap state Keimer et al. (2015). In contrast, the overdoped region is free from competing phases, but its normal state ($T>T_{c}$) shows “strange metal” (SM) behaviour Phillips et al. (2023) where the resistivity, $\rho$, is proportional to temperature, $T$, with a large proportionally constant.
The resistivity of a metal is usually interpreted using the Drude model in which $\rho\propto\tau_{\text{tr}}^{-1}$, where $\tau_{\text{tr}}^{-1}$ is the relaxation rate of the electrons transporting current. SM behaviour occurs in many classes of material Hartnoll and Mackenzie (2022) where it has been found that $\tau_{\text{tr}}^{-1}$ is approximately equal to $k_{B}T/\hbar$, the Planckian dissipative rate Zaanen (2004). This is in stark contrast to the $\rho\propto T^{2}$ and $\tau_{\text{tr}}^{-1}\propto T^{2}$ expected to describe metals in Fermi liquid theory at low temperature. The SM Hartnoll and Mackenzie (2022) in resistivity can be accompanied by an anomalous specific heat contribution $C\propto T\ln(1/T)$ in contrast to the liquid Fermi form $C\propto T$.
The existence of SM behaviour over a wide $T$ range in cuprates is challenging for theory to explain Patel et al. (2023). The picture in simple metals is based on the current-carrying fermion quasiparticles near the Fermi surface scattering off bosons (e.g. phonons). Phonons can only give a linear behaviour at high temperatures $T\gtrsim\theta_{\text{Debye}}$ where all phonon modes are excited. Electron-electron scattering yields $\rho\propto T^{2}$ because only electrons within $k_{B}T$ of the Fermi surface scatter. In addition, the Planckian dissipation mentioned above implies that quasiparticles lose coherence in the shortest time allowed by quantum mechanics and that the quasiparticle concept breaks down.
@@
Here we show that nearly-critical low-energy collective spin fluctuations are central to the strange metal behaviour in the cuprates and exist across the superconducting dome. We investigate the overdoped region of a cuprate superconductor where the SM behaviour is clearest Cooper et al. (2009). The material we have chosen is La2-xSrxCuO4 (LSCO). This system can be hole-doped over a wide range through the overdoped region of the phase diagram (Fig. 1A) and has a relatively low $T_{c}$ so that the SM normal state is present over approximately two orders of magnitude in temperature.
@@
Figure 1: Spin fluctuations in La2-xSrxCuO4 (x=0.22). (A) Schematic phase diagram of LSCO showing the possible extent of the critical spin fluctuations observed here. The dashed line shows the extent of spin freezing for $B=80$ T Frachet et al. (2020) where superconductivity is fully suppressed and the magenta region for $B=0$ from Ref. Kofu et al. (2009). (B) Schematic of spin excitations in LSCO($x=0.22$) based on (Refs. Lipscombe et al. (2007); Robarts et al. (2019)) as represented by contours of the imaginary part of the susceptibility $\chi^{\prime\prime}(\mathbf{Q},\omega)$. Reciprocal space is labelled as $\mathbf{Q}=H\mathbf{a}^{\star}+K\mathbf{b}^{\star}+L\mathbf{c}^{\star}\equiv(H,K)$. The gray region shows the broad high-energy spin excitations Lipscombe et al. (2007); Robarts et al. (2019). Pink regions are the low-energy excitations studied here. (C-F) Slices of $\chi^{\prime\prime}(\mathbf{Q},\omega)$ for $\hbar\omega=4$ meV, $L\in[-1,1]$ and various temperatures (see Methods for details). The white rectangle in (D) shows the region of integration used to produce the points in 1-D cuts such as those in Fig. 2A, with integration along (1,1). Dashed lines represent the path of the 1-D cuts.
@@
Figure 2: The temperature dependence of the low-energy spin fluctuations in La2-xSrxCuO4 (x=0.22). (A) The scattering intensity versus wavevector $\mathbf{Q}$ for different temperatures (bottom to top: 26, 40, 80, 155, 300 K) for energy transfers $\hbar\omega=4,8$ meV. The cuts are through the incommensurate $\mathbf{Q}_{\delta}$ positions as shown by the dashed lines in Fig. 1D. (B) $E-\mathbf{Q}$ map of the magnetic response function $\chi^{\prime\prime}(\mathbf{Q},\omega)$ for $T=26,80$ K showing its evolution with temperature. The trajectory of $\mathbf{Q}$ is the same as that in (A).
@@
Figure 3: Scaling behaviour of the dynamic magnetic susceptibility and correlation length in La2-xSrxCuO4 (A) Temperature dependence of $\chi^{\prime\prime}(\mathbf{Q},\omega)$ at ordering wavevector $\mathbf{Q}_{\delta}$. Lines are fits to Eqn. 3. Open and closed symbols are two separate experiments. Inset shows the $T$-dependence of the spin relaxation rate $\Gamma_{\delta}$ and real part of the susceptibility $\chi^{\prime}(\mathbf{Q}_{\delta})$ obtained when data are modelled with Eqns. 3-4. (B) $\omega/T$-scaling plot of the $x=0.22$ data in (A) using Eqn. 2 and the procedure described in Methods. The inset shows how the quality of the collapse varies with exponent $\alpha$. (C) Scaling plot for underdoped LSCO $x=0.14$ data from Ref. Aeppli et al. (1997). This data shows a similar collapse to the $x=0.22$ data in (B). (D) The width in $\mathbf{Q}$ of $\chi^{\prime\prime}(\mathbf{Q,\omega})$ is denoted by the inverse dynamic correlation length $\kappa(\omega)$. Here we plot $\kappa(\omega)$ versus temperature for $\hbar\omega=1.25,2.25,4,8$ meV. (E) Data in (D) can be collapsed onto a single trend line with suitable choice of $z$ and $\omega/T$-scaling. (F) Data in (D) can be scaled onto a single line using (7) allowing the dynamic critical exponent $z$ to be determined. Dashed line is Supplementary Equation 15.
Results
Spin fluctuations in cuprates
@@ Results
We used inelastic neutron scattering to map out the $\mathbf{Q}$-$\omega$ dependence of the low-energy spin fluctuations in single crystals of LSCO(x=0.22=0.22) in the normal state for temperatures $T_{c}\leq T\leq 300$ K. Measurements were carried out using the LET spectrometer of the ISIS Pulsed Neutron and Muon Source (See Methods). Fig. 1C-F show constant energy maps of the dynamic susceptibility $\chi^{\prime\prime}(\mathbf{Q},\omega)$ at $\hbar\omega=4$ meV. For $T=26$ K, the four-peaks around $\mathbf{Q}=(1/2,1/2)$ can be clearly seen in Fig. 1C. Cuts through the peaks along the $\mathbf{Q}=(1/2-\delta/2+\xi,1/2-\delta/2+\xi)$ line (along the centre of dashed lines in Fig. 1D) are shown in Fig. 2A. As the temperature increases, the peaks can be seen to weaken and broaden in $\mathbf{Q}$. The measured intensity $S(\mathbf{Q},\omega)$ can be converted to the the magnetic response function $\chi^{\prime\prime}(\mathbf{Q},\omega)$ using the fluctuation dissipation theorem $S(\mathbf{Q},\omega)=(1/\pi)\chi^{\prime\prime}(\mathbf{Q},\omega)[1-\exp(-\hbar\omega/k_{B}T)]^{-1}$ (See Methods), this is shown in Fig. 2B and Fig. 3A. Our data (see Fig. 3A) show that the dynamic spin susceptibility near $\mathbf{Q}_{\delta}$ is strongly temperature dependent for this overdoped cuprate. The behaviour is reminiscent of the approach to a quantum critical point as $T\rightarrow 0$.
.1 Modelling the low-energy spin fluctuations
At a quantum critical point, the dynamical susceptibility is expected to have the following scaling form Sachdev and Ye (1992)
@@
where $\xi\propto T^{-\frac{1}{z}}$ is a correlation length, $z$ is the dynamic critical exponent, $\mathbf{q}=\mathbf{Q}-\mathbf{Q}_{\delta}$ is the wavevector relative to the ordering wavevector, $q=|\mathbf{q}|$ and $\Phi$ is a universal complex function of both arguments. For $q=0$, the imaginary part of the dynamical susceptibility (measured here) has the following scaling form (See Supplementary Methods):
@@
where $\alpha=\gamma/\nu z$, $\phi_{1}(x)=\Im[\Phi(0,x)]$ is a scaling function and $\gamma$ and $\nu$ are the susceptibility and correlation function temperature exponents respectively. This is an example of $\omega/T$-scaling.
In Fig. 3B we show that our data can be collapsed onto a single trend with a suitable choice of $\alpha$. An analogous collapse is observed in heavy fermion systems Aronson et al. (1995a); Schröder et al. (2000). In our case, $\phi_{1}(x)$ can be approximated by a simple Lorentzian $\phi_{1}(x)\propto ax/(a^{2}+x^{2})$ with $a=2.9\pm 0.3$ and $\alpha=0.32\pm 0.10$. This small value of $\alpha$ excludes the familiar paramagnon theory which has $\alpha=1$ Millis et al. (1990); Millis (1993) and does not show an extended regime of criticality. Our data are consistent with the measured susceptibility having a low-frequency component that varies as
@@
where the real part of the susceptibility $\chi^{\prime}(\mathbf{Q}_{\delta})\propto T^{-\alpha}$ and the spin relaxation rate varies as
@@
as shown in Fig. 3A(inset). Thus our data are consistent with $\Gamma_{\delta}$ tracking the Planckian relaxation rate $\tau^{-1}_{\text{tr}}\propto T$ as observed in transport measurements suggesting that they are related.
More information about the criticality of the spin fluctuations can be gained from the $\mathbf{Q}$-dependence of $\chi^{\prime\prime}(\mathbf{Q},\omega)$. We fitted our data using a form previously used Aeppli et al. (1997); Zhu et al. (2023) to describe the low-energy response of the cuprates:
$\chi^{\prime\prime}(\mathbf{Q},\omega)$ $=\chi^{\prime\prime}(\mathbf{Q}_{\delta},\omega)\frac{1}{[1+\kappa^{-2}(\omega)R(\mathbf{Q})]^{2}},$ (5)
where $R(\mathbf{Q})$ is a function with zeros at the $\mathbf{Q}_{\delta}$ positions and $R(\mathbf{Q})=|\mathbf{q}|^{2}$ near $\mathbf{Q}=\mathbf{Q}_{\delta}$ [see Supplementary Equation (11)]. Eqn. 5 has peaks at the four $\mathbf{Q}_{\delta}$ positions and $\kappa(\omega)$ is a measure of the peak width. Example fits to Eqn. 5 are shown in Fig. 2 and the resulting values of $\kappa(\omega)$ are shown in Fig. 3D,E,F. Eqns. 1 and 5 taken together (See Supplementary Methods) imply an additional $\kappa$-scaling
$\kappa(\omega)=T^{\tfrac{1}{z}}\phi_{2}\left(\frac{\hbar\omega}{k_{B}T}\right)$ (6)
and $\kappa(0)\propto T^{\frac{1}{z}}$.
In Fig. 3D we see that $\kappa(\omega)$ increases with $\omega$ and $T$. We posit that the effects of $\omega$ and $T$ add as
@@
where $r$ is the ratio of the temperature and energy contributions which we set equal to one. This form satisfies Eqn. 6. We find that Eqn. 7 can be used to collapse the experimentally determined $\kappa(\omega,T)$ onto a single solid line as shown in Fig. 3E,F with $z=1.83\pm 0.35$. In the Supplementary Methods we show that our data are also consistent with the system being tuned slightly away from quantum criticality (dashed line in Fig. 3F) within the uncertainty of the experiment.